# An extension of Birkhoff’s representation theorem to locally-finite distributive lattices

### 06 Mar 2026 v1

@@ 2. Representation of non-finitary lattices
[Birk1937a][Birk1967a]*§III.3 p. 58[Stan2012a]*Prop. 3.4.3 If $\mathcal{L}$ is finitary then every element $p\in\mathcal{F}$ is ${x}\mathclose{\uparrow}$} for some $x\in\mathcal{L}$.
@@
If $\mathcal{L}$ is finitary then: (1) $\mathcal{P}$ has no sur-primes, so $\mathcal{P}$ is (isomorphic to) the poset of join-irreducible elements of $\mathcal{L}$. (2) The image of $\phi$ is the set of finite elements of $\mathcal{I}$, so $\phi$ is a lattice isomorphism of $\mathcal{L}$ to that set.
Application to locally-finite lattices
In this section, we additionally assume that $\mathcal{L}$ is locally-finite (but not necessarily finitary).
@@
If $x\in\mathcal{L}$ is join-irreducible, then $x$ covers exactly one element of $\mathcal{L}$.
@@
*** may no longer be needed
Definition 31.
For any $x\in\mathcal{L}$ that is join-irreducible, we define $x^{-}$ (relative to $\mathcal{L}$) to be the unique element that $x$ covers.
@@
Dually, we define the parallel operations on lattice ideals:
@@
By the definition of $\phi$, it is clear that $\phi(x)\subset\phi(y)$.
By cor. 20, there exists a $p\in\mathcal{P}$ for which $y\in p$ and $x\not\in p$. Thus $p\in\phi(y)\setminus\phi(x)$.
@@
Given a covering $x\lessdot y$, we define the separator of $x$ and $y$, $x\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y$ to be unique element of $\mathcal{P}$ that is in $\phi(y)\setminus\phi(x)$. (lem. 36
@@
Regarding $\phi(y)\subset\phi(x)\sqcup\{p\}$: For any $q\in\phi(y)$, $q\leq{{y}\mathclose{\uparrow}}={{x}\mathclose{\uparrow}}\vee p$. Since $\mathcal{P}$ is distributive and $q$ is join-irreducible in $\mathcal{P}$, $q\leq{{x}\mathclose{\uparrow}}$ or $q\leq p$. That is equivalent to $q\in\phi(x)$ or $q\leq p$. Since $\phi(x)\sqcup\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\in\phi(x)\sqcup\{p\}$.
By lem. 32, $x<y$, and since $\phi(y)$ has exactly one more element than $\phi(x)$, lem. 38 shows that $x\lessdot y$. ∎
@@
Suppose $x\in\mathcal{L}$ and $p\in\mathcal{P}$ such that $p\in\phi(x)$ and $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$. Then $x\gtrdot x\mathord{\Downarrow}(\mathcal{L}\setminus p)$ and $\phi(x\mathord{\Downarrow}p)=\phi(x)\setminus\{p\}$.
@@
The proof is dual to the proof of lem. 39 but that is concealed by our notation, since $\phi$ is not self-dual.
@@
Regarding $\phi(x)\setminus\{p\}\subset\phi(y)$: We prove this by the contrapositive: For any $q\not\in\phi(y)$, $y\not\in q$. Define $q^{\prime}=\mathcal{L}\setminus q$, so $q^{\prime}$ is a lattice ideal and $y\in q^{\prime}$. Since $q$ is join-irreducible in $\mathcal{F}$, $q^{\prime}$ is meet-irreducible in $\mathcal{F}^{\prime}$. Then computing in $\mathcal{F}^{\prime}$, $q^{\prime}\geq{{y}\mathclose{\downarrow}}={{x}\mathclose{\downarrow}}\wedge p^{\prime}$. Since $q^{\prime}$ is meet-irreducible and $\mathcal{F}^{\prime}$ is distributive, $q^{\prime}\geq{{x}\mathclose{\downarrow}}$ or $q^{\prime}\geq p^{\prime}$. That is equivalent to $x\in q^{\prime}$ or $q^{\prime}\geq p^{\prime}$, which is equivalent to $x\not\in q$ or $q\geq p$, which is further equivalent to $q\not\in\phi(x)$ or $q\geq p$. Since $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\not\in\phi(x)\setminus\{p\}$.
@@
By th. 26, $\phi$ is a lattice embedding of $\mathcal{L}$ into $\mathcal{I}$, so what remains to be shown is that the image of $\phi$ is $D_{\phi(x)}$.
@@
Thus, if we look at $\mathcal{I}$ as a graph on its elements with two elements being connected iff one covers the other, $\mathcal{L}$ is isomorphic to one connected component of $\mathcal{I}$.
In the example $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ in fig. 6, $\mathcal{L}$ is isomorphic to the central connected component and there are eight other connected components. In this case, if we choose any of the eight other connected components as $\mathcal{L}^{\prime}$ and constructed the corresponding $\mathcal{I}^{\prime}$, it would be different from this $\mathcal{I}$.
However, other examples behave differently. If we choose $\mathcal{L}=\mathbb{B}_{\textup{fin}}$, the lattice of finite subsets of $\mathbb{N}$, then the corresponding $\mathcal{P}$ is isomorphic to $\mathbb{N}$, with each prime filter being all the finite sets containing some particular integer. The corresponding $\mathcal{I}$ is isomorphic to $\mathbb{B}_{\mathbb{N}}$, the lattice of all subsets of $\mathbb{N}$. Then $\mathcal{L}$ is isomorphic to the bottom connected component of $\mathcal{I}$, containing its $\hat{0}$.
The top connected component of $\mathcal{I}$ is isomorphic to $\mathbb{B}_{\textup{cofin}}$, the lattice of all cofinite subsets of $\mathbb{N}$, which is the lattice dual of $\mathbb{B}_{\textup{fin}}$. All other connected components are isomorphic to $\mathbb{B}_{\textup{fin}}\times\mathbb{B}_{\textup{cofin}}$. And since all components are countable but $\mathbb{B}_{\mathbb{N}}$ is uncountable, there are an uncountable number of connected components.
If we chose $\mathcal{L}$ to be any one of these components, then $\mathcal{P}$ would still be isomorphic to $\mathbb{N}$ and $\mathcal{I}$ would be the same. In this situation, for every component and every prime filter, there is some covering in the component whose separator is the prime filter.
*** Need to characterize when all prime filters are principal.
@@
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ and $x^{+}/x^{-}\searrow z^{+}/z^{-}$, then $w^{-}\lessdot w^{+}$, $y^{+}/y^{-}\searrow w^{+}/w^{-}$, and $z^{+}/z^{-}\searrow w^{+}/w^{-}$, where $w^{-}=y^{-}\wedge z^{-}$ and $w^{+}=y^{+}\wedge z^{+}$.
@@
The diamond isomorphism theorem shows that $w^{-}\lessdot w^{+}$. *** ∎
Downward transposes preserve separators:
@@
Lemma 48.
@@
    if $p$ is principal and $p={{x^{+}}\mathclose{\uparrow}}$, then there is no $x^{+}/x^{-}\searrow y^{+}/y^{-}$,
@@
    otherwise, there exists a $x^{+}/x^{-}\searrow y^{+}/y^{-}$.
@@
    if $p={{y}\mathclose{\uparrow}}$, then any chain of downward transposes of $x^{-}\lessdot x^{+}$ can be extended by appending $y^{-}\lessdot y$.
@@
    if $p$ is not principal, then all chains of downward transposes of $x^{-}\lessdot x^{+}$ can be extended indefinitely.

### 15 Mar 2026 v2

@@ 3. Application to finitary lattices
If $\mathcal{L}$ is finitary then every element $p\in\mathcal{F}$ is ${x}\mathclose{\uparrow}$} for some $x\in\mathcal{L}$.
@@
[Stan2012a]*Prop. 3.4.3 If $\mathcal{L}$ is finitary then: (1) $\mathcal{P}$ contains no secondary primes, so $\mathcal{P}$ is (isomorphic to) the poset of join-irreducible elements of $\mathcal{L}$. (2) Every value of $\phi$ is a finite set in $\mathcal{I}$. (3) For every finite $I\in\mathcal{I}$, $\phi(\bigvee I)=I$. (4) $\phi$ is a lattice isomorphism between $\mathcal{L}$ and the set of finite elements of $\mathcal{I}$.
4. Application to locally-finite lattices
In this section, we apply the general theory to lattices that are locally-finite but not necessarily finitary. Henceforth we assume that $\mathcal{L}$ is locally-finite.
@@ 4. Application to locally-finite lattices
If $x\in\mathcal{L}$ is join-irreducible, then $x$ covers exactly one element of $\mathcal{L}$. Dually, if $x\in\mathcal{L}$ is meet-irreducible, then $x$ is covered by exactly one element of $\mathcal{L}$. (The converses of these statements are trivial.)
@@
We define the dual operation on lattice ideals:
@@
By def. 25, it is clear that $\phi(x)\subset\phi(y)$.
By cor. 21, there exists a $p\in\mathcal{P}$ for which $y\in p$ and $x\not\in p$. Thus $p\in\phi(y)\setminus\phi(x)$.
@@
Given a covering $x\lessdot y$, we define the separator of $x$ and $y$, $x\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y$ to be the unique element of $\phi(y)\setminus\phi(x)$. (Thus $x\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y\in\mathcal{P}$.)
@@
Regarding $\phi(y)\subset\phi(x)\sqcup\{p\}$: For any $q\in\phi(y)$, $q\leq{{y}\mathclose{\uparrow}}={{x}\mathclose{\uparrow}}\vee p$. Since $\mathcal{P}$ is distributive and $q$ is join-irreducible in $\mathcal{P}$ by lem. 23, $q\leq{{x}\mathclose{\uparrow}}$ or $q\leq p$. That is equivalent to $q\in\phi(x)$ or $q\leq p$. Since $\phi(x)\sqcup\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\in\phi(x)\sqcup\{p\}$. Thus $\phi(y)\subset\phi(x)\sqcup\{p\}$.
Together, these show $\phi(y)=\phi(x)\sqcup\{p\}$. By lem. 32, $x<y$, and since $\phi(y)$ has exactly one more element than $\phi(x)$, lem. 38 shows that $x\lessdot y$. ∎
@@
Suppose $x\in\mathcal{L}$ and $p\in\mathcal{P}$ such that $p\in\phi(x)$ and $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$. Then defining $p^{\prime}=\mathcal{L}\setminus p$, $x\gtrdot x\mathord{\Downarrow}p^{\prime}$ and $\phi(x\mathord{\Downarrow}p^{\prime})=\phi(x)\setminus\{p\}$.
@@
(This proof is dual to the proof of lem. 39 but that is concealed by our notation, since $\phi$ is not self-dual.)
@@
Regarding $\phi(x)\setminus\{p\}\subset\phi(y)$: We prove this by the contrapositive: For any $q\not\in\phi(y)$, $y\not\in q$. Define $q^{\prime}=\mathcal{L}\setminus q$, so $q^{\prime}$ is a lattice ideal and $y\in q^{\prime}$. Since $q$ is join-irreducible in $\mathcal{F}$ by lem. 23, $q^{\prime}$ is meet-irreducible in $\mathcal{F}^{\prime}$. Then computing in $\mathcal{F}^{\prime}$, $q^{\prime}\geq{{y}\mathclose{\downarrow}}={{x}\mathclose{\downarrow}}\wedge p^{\prime}$. Since $q^{\prime}$ is meet-irreducible and $\mathcal{F}^{\prime}$ is distributive, $q^{\prime}\geq{{x}\mathclose{\downarrow}}$ or $q^{\prime}\geq p^{\prime}$. That is equivalent to $x\in q^{\prime}$ or $q^{\prime}\geq p^{\prime}$, which is equivalent to $x\not\in q$ or $q\geq p$, which is further equivalent to $q\not\in\phi(x)$ or $q\geq p$. Since $\phi(x)\setminus\{p\}$ is an order ideal of $\mathcal{P}$ by hypothesis, either case implies $q\not\in\phi(x)\setminus\{p\}$.
@@
By th. 27, $\phi$ is a lattice embedding of $\mathcal{L}$ into $\mathcal{I}$, so what remains to be shown is that the image of $\phi$ is $D_{\phi(x)}$.
@@
Thus, if we look at $\mathcal{I}$ as a graph whose vertices are the elements of $\mathcal{I}$ with two elements being connected iff one covers the other, $\mathcal{L}$ is isomorphic to one connected component of $\mathcal{I}$.
In the example $\mathcal{L}=\mathbb{Z}\times\mathbb{Z}$ in fig. 6, $\mathcal{L}$ is isomorphic to the central connected component, and there are eight other connected components. In this case, if we chose any of the eight other connected components as $\mathcal{L}_{1}$, the corresponding $\mathcal{P}_{1}$ would be a proper subset of $\mathcal{P}$, and $\mathcal{I}_{1}$ would be a proper subset of this $\mathcal{I}$.
However, other examples behave differently. If we choose $\mathcal{L}=\mathbb{B}_{\textup{fin}}$, the lattice of finite subsets of $\mathbb{N}$, then the corresponding $\mathcal{P}$ is an antichain with $\aleph_{0}$ elements, with each prime filter being all the finite sets containing some particular integer and each prime filter is a principal prime. The corresponding $\mathcal{I}$ is isomorphic to $\mathbb{B}_{\mathbb{N}}$, the lattice of all subsets of $\mathbb{N}$. Then $\mathcal{L}$ is isomorphic to the bottom connected component of $\mathcal{I}$, containing its $\hat{0}$$\emptyset$.
The top connected component of $\mathcal{I}$ contains its $\hat{1}=\mathcal{P}$ and is isomorphic to $\mathbb{B}_{\textup{cofin}}$, the lattice of all cofinite subsets of $\mathbb{N}$, which is the lattice dual of $\mathbb{B}_{\textup{fin}}$. All other connected components are isomorphic to $\mathbb{B}_{\textup{fin}}\times\mathbb{B}_{\textup{cofin}}$. And since all components are countable but $\mathbb{B}_{\mathbb{N}}$ is uncountable, there are an uncountable number of connected components of $\mathcal{I}$.
If we chose $\mathcal{L}_{1}$ to be any one of these components other than $\mathbb{B}_{\textup{fin}}$, then $\mathcal{P}_{1}$ would be an antichain with $\aleph_{0}$ elements, but its elements would all be secondary primes. The corresponding $\mathcal{I}_{1}$ would be isomorphic to $\mathcal{I}$.
5. Criterion for secondary primes
Given a locally-finite lattice, if all of its primes are principal, they correspond to the join-irreducible elements of $\mathcal{L}$, and so the structure of $\mathcal{P}$ is relatively easy to determine. But if a prime is secondary, it has no single corresponding element of $\mathcal{L}$ and so determining the structure of $\mathcal{P}$ is more difficult. We will now show how to distinguish these two cases.
@@ 5. Criterion for secondary primes
If $x^{+}/x^{-}\searrow y^{+}/y^{-}$ and $x^{+}/x^{-}\searrow z^{+}/z^{-}$, then defining $w^{-}=y^{-}\wedge z^{-}$ and $^{+}=y^{+}\wedge z^{+}, ww$$^{-}\lessdot w^{+}$, $y^{+}/y^{-}\searrow w^{+}/w^{-}$, and $z^{+}/z^{-}\searrow w^{+}/w^{-}$.
@@
By hypothesis, $y^{-}=y^{+}\wedge x^{-}$, $z^{-}=z^{+}\wedge x^{-}$, $x^{+}=x^{-}\vee z^{+}$, and $y^{+}\leq x^{+}$. Then $y^{-}\wedge w^{+}=y^{-}\wedge(y^{+}\wedge z^{+})=(y^{+}\wedge x^{-})\wedge(y^{+}\wedge z^{+})=(y^{+}\wedge x^{-})\wedge(z^{+}\wedge x^{-})=y^{-}\wedge z^{-}=w^{-}$ and $y^{-}\vee w^{+}=(y^{+}\wedge x^{-})\vee(y^{+}\wedge z^{+})=y^{+}\wedge(x^{-}\vee z^{+})=y^{+}\wedge x^{+}=y^{+}$. Together, these show $y^{+}/y^{-}\searrow w^{+}/w^{-}$. Similarly, we show $z^{+}/z^{-}\searrow w^{+}/w^{-}$. Applying the diamond transposition theorem to either downward transpose shows that $w^{-}\lessdot w^{+}$. ∎
Downward transposes preserve separators (def. 37):
@@
We know that $\phi(x^{+})\setminus\phi(x^{-})$ and $\phi(y^{+})\setminus\phi(y^{-})$ each contain exactly one element. Define $p=x^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}x^{-}=\phi(x^{+})\setminus\phi(x^{-})$, so $p\not\in\phi(x^{-})$, $x^{-}\not\in p$, and $x^{-}\vee y^{+}=x^{+}\in p$. Since $p$ is a prime filter, either $x^{-}\in p$ or $y^{+}\in p$. But by hypothesis, $x^{-}\not\in p$, so $y^{+}\in p$. Since $x^{-}\not\in p$ by hypothesis and $y^{-}\leq x^{-}$, $y^{-}\not\in p$. Thus $y\in\phi(y^{+})\setminus\phi(y^{-})$, so $p=y^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y^{-}$. ∎
Theorem 48.
@@
    if $p$ is principal and $p={{x^{+}}\mathclose{\uparrow}}$, then there is no $x^{+}/x^{-}\searrow y^{+}/y^{-}$ with $x^{+}>y^{+}$.
@@
    otherwise, there exists a $y^{-}\lessdot y^{+}$ with $x^{+}>y^{+}$ and $x^{+}/x^{-}\searrow y^{+}/y^{-}$.
Proof.
Regarding (1): If there was such a $y^{-}\lessdot y^{+}$, then by lem. 47, $p=y^{+}\mathord{\mathchoice{/\mkern-5.0mu/}{/\mkern-4.0mu/}{/\mkern-5.0mu/}{/\mkern-5.0mu/}}y^{-}$ and $y^{+}\in p$. But that would mean that $y^{+}\geq x^{+}$, but by hypothesis $x^{+}>y^{+}$.
Regarding (2): If $p$ is not principal, or $p\neq{{x^{+}}\mathclose{\uparrow}}$, then there exists a $y^{+}\lessdot x^{+}$, and necessarily $y^{+}\neq x^{-}$. Define $y^{-}=y^{+}\wedge x^{-}$. Then by the diamond isomorphism theorem, $y^{-}\lessdot y^{+}$ and $x^{+}/x^{-}\searrow y^{+}/y^{-}$. ∎
Definition 49.
We define a chain of downward transposes to be a sequence $x^{-}_{1}\lessdot x^{+}_{1},x^{-}_{2}\lessdot x^{+}_{2},x^{-}_{3}\lessdot x^{+}_{3},\ldots,x^{-}_{n}\lessdot x^{+}_{n}$ with $x^{+}_{1}>x^{+}_{2}>x^{+}_{3}>\cdots>x^{+}_{n}$ (or equivalently $x^{-}_{1}>x^{-}_{2}>x^{-}_{3}>\cdots>x^{-}_{n}$) and $x_{1}^{+}/x_{1}^{-}\searrow x_{2}^{+}/x_{2}^{-}$, $x_{2}^{+}/x_{2}^{-}\searrow x_{3}^{+}/x_{3}^{-}$, $\ldots$, $x_{n-1}^{+}/x_{n-1}^{-}\searrow x_{n}^{+}/x_{n}^{-}$.
@@
    if $p$ is principal, then any chain of downward transposes of starting with $x^{-}\lessdot x^{+}$ can only be extended to some finite maximum length.
@@
    if $p$ is secondary, then all chains of downward transposes starting with $x^{-}\lessdot x^{+}$ can be extended indefinitely.
Cor. 50 can be used to discern which prime filters are primary and which are secondary. For example, in fig. 7(a) all chains of downward transposes can be extended indefinitely, so all primes are secondary.
But in fig. 7(b) all chains of downward transposes must end, showing all primes are principal. The upper elements of the coverings on which chains are forced to end are the join-irreducible elements of $\mathcal{L}$, and collectively they are are isomorphic to $\mathcal{P}$. $\mathcal{I}$ has three connected components, one is $\emptyset$, one is $\mathcal{P}$, and one is the remaining ideals of $\mathcal{P}$, all of which are infinite and have complements that are infinite. The third connected component is isomorphic to this lattice.
(a) The lattice $\mathbb{Z}\times\mathbb{Z}$ with a chain of downward transposes
(b) The lattice 44-SkewStrip with several chains of downward transposes
Figure 7.
6. Unique connected component of ideals
The examples we have been using, $\mathbb{Z}\times\mathbb{Z}$ and $\mathbb{B}_{\textup{fin}}$, share the property that their lattice of ideals $\mathcal{I}$ has multiple connected components, one of which is isomorphic the the lattice that generates it. However, our long-term goal is characterizing lattices which are useful in constructing RSK algorithms. Examples of these lattices are $t$-SkewStrip in fig. 3 and the lattices listed in [Wor2026b]*sec. 2.2. All of those lattices have the property that that their lattice of ideals $\mathcal{I}$ has at most three components:
  • one component that is isomorphic to the generating lattice $\mathcal{L}$;
  • if $\mathcal{L}$ has no $\hat{1}$, one component containing only $\mathcal{P}$; and
  • if $\mathcal{L}$ has no $\hat{0}$, one component containing only $\emptyset$.
Looking at these examples suggests:
Conjecture 51.
A locally-finite distributive lattice has a lattice of ideals of prime filters that contains at most three connected components (one isomorphic to the lattice, perhaps one containing the empty ideal, and perhaps one containing the ideal of all primes) iff the lattice has finite width (contains no infinite antichains).